OpenStudy (anonymous):

find the slope of the line that passes through the pair of points (-4,2), (-2,5)

OpenStudy (anonymous):

To find the slope, consider the change in x and the change in y. Over a change in x of 2 (as -4 changes to -2), y changes by 3 (changes from 2 to 5). So the slope of the line is the change in y over the change in x: 3/2. You can use a formula to find this: \[slope = \frac{ \Delta y }{ \Delta x } = \frac{ y_2 - y_1 }{ x_2 - x_1 } = \frac{ 5 - 2 }{ -2 - (-4) } = \frac{ 3 }{ 2 }\]